OpenStudy (anonymous):

A recipe for 12 waffles calls for 1 ½ cups of milk, 2 ¼ cups of flour and 1 1/3 cups of other ingredients. How many cups of flour, milk, and other ingredients are needed to make 48 waffles is the answer 3 cups milk, 4 1/2 cups flour, and 2 2/3 cups other?

OpenStudy (callisto):

48 waffles = 4 (12 waffles) So just multiply all ingredients by 4 Milk = 1.5 *4 = 6 cups flour = 2.25 *4 = 9 cups and so on, try again yourself :)

OpenStudy (callisto):

for other ingredients , it is = 4 (4/3) = 16/3 = 5 1/3

OpenStudy (anonymous):

thank you so much, i got it right i just had a blonde moment for a second!

OpenStudy (callisto):

welcome